OpenStudy (anonymous):

What is the period of the sinusoid given by ? y= -2sin(2pi/7x)

OpenStudy (anonymous):

For a sinusoid of the form \(A\sin (B(x-C))+D\) (or \(\cos\) in place of \(\sin\)), the period is given by \(\dfrac{2\pi}{B}\). Your function has \(B=\dfrac{2\pi}7\).